# Visual Whole-Body Control for Legged Loco-Manipulation

### v3

@@ Abstract
We study the problem of mobile manipulation using legged robots equipped with an arm, namely legged loco-manipulation. The robot legs, while usually utilized for mobility, offer an opportunity to amplify the manipulation capabilities by conducting whole-body control. That is, the robot can control the legs and the arm at the same time to extend its workspace. We propose a framework that can conduct the whole-body control autonomously with visual observations. Our approach, namely Visual Whole-Body Control (VBC), is composed of a low-level policy using all degrees of freedom to track the end-effector manipulator position and a high-level policy proposing the end-effector position based on visual inputs. We train both levels of policies in simulation and perform Sim2Real transfer for real robot deployment. We perform extensive experiments and show significant improvements over baselines in picking up diverse objects in different configurations (heights, locations, orientations) and environments.
@@ I Introduction
An important aspect of loco-manipulation is the opportunity to conduct whole-body control using all degrees of freedom of the robot. While the legs are typically utilized for mobility, there is a large potential for controlling all the joints of the legs and arms together to amplify manipulation capabilities. We humans do this, too: We use our legs to stand up to reach a book placed on a shelf; We bend our legs to tie our shoes. Fig. 1 shows the tasks our robot can do when controlling the whole-body joints together. For example, our robot can go to the side of the road and pick up the trash on the grass. This will not be achievable without bending the front legs of the robot. The arm alone will not be long enough to reach the ground.
@@ II Related Work
Legged Locomotion. For decades, legged locomotion has been an essential topic in the field of robotics and has gained much success. Classical methods solve the legged locomotion task through model-based control to define controllers in advance. For example, Miura and Shimoyama [50], Sreenath et al. [61] utilized traditional methods to enable the walking of biped robots. MIT cheetah 3 [7], ANYmal-a [31] by ETH, and Atlas [19] by Boston Dynamics are all perfect examples of robots designed with a robust walking controller. Based on these developed robots, researchers have been able to develop advanced controllers for even more agile locomotion tasks, such as jumping or walking over varying obstacles [53, 51, 26, 52, 3, 38]. To make a robust controller that can adapt to varying environments, researchers integrate some sensor data such as elevation maps [18, 40, 35, 6, 17] and state estimators [8, 66, 10, 74, 65]. Although classical methods have greatly succeeded in legged locomotion, generalizability is always a problem for pre-designed controllers, and acquiring these controllers requires careful and tedious engineering. With the fast development of deep neural networks, learning-based end-to-end methods have become a simple way to gain robust controllers with strong generalization abilities. Learning-based methods often train a robust controller in physics simulators such as PyBullet [14] and Isaac Gym [47] and then perform sim-to-real transfer to a real robot. For example, Rudin et al. [58] uses Isaac Gym with GPU to enable a quadruped robot to walk in only several minutes, Kumar et al. [36, 37] proposed the RMA algorithm to make the sim-to-real transfer process easier, and more robust, Lee et al. [41], Miki et al. [49], Agarwal et al. [1], Yang et al. [73], Yu et al. [80], Cheng et al. [13], Zhuang et al. [82], Fu et al. [21], Margolis and Agrawal [48], Li et al. [44] enable the legged robots to walk or run on varying terrains, and Fuchioka et al. [24], Escontrela et al. [16], Peng et al. [54], Yang et al. [71], Wang et al. [64], Li et al. [43], Kumar et al. [39] combine RL with imitation learning to produce naturalistic and robust controllers for legged robots. Kim et al. [33] map human motions to a quadruped robot to complete a wide range of locomotion and manipulation tasks.
@@
Legged Loco-Manipulation A set of works has exploited legged loco-manipulation to take advantage of legged robots. Bellicoso et al. [5] designs an online motion planning framework, together with a whole-body controller based on a hierarchical optimization algorithm to achieve loco-manipulation, Fu et al. [22] proposed an effective method to train a whole-body controller end-to-end through DRL. Ma et al. [46] combined model-based manipulator control and learning-based locomotion, but only showed the response to external push instead of achieving manipulation tasks. Notably, these works all require human involvement and teleoperation to execute and achieve tasks, rather than operating autonomously. Zhang et al. [81], Yokoyama et al. [79], Arcari et al. [4] all performed an autonomous mobile manipulation system on quadruped robots while they utilized the default model-based controller provided through the robot APIs (only control robot locomotion in a static height) and only focused on the manipulation part without any whole-body behaviors to adapt to different heights of objects. Ferrolho et al. [20] proposed a trajectory optimization framework based on a robustness metric for solving complex loco-manipulation tasks, but the task is assigned and the target positions are calculated via a motion capture system, thus only works limitedly. Sleiman et al. [60] realized a bilevel search strategy for motion planning, by incorporating domain-specific rules, combing trajectory optimization, informed graph search, and sampling-based planning, enabling robots to perform complex tasks. They require the user to provide a dense description of the scene to initialize the solver. Zimmermann et al. [83] also utilized trajectory optimization to enable a legged robot to walk and pick up objects. However, they only show limited surroundings and simple objects. Besides these works with arm-based configuration, Wolfslag et al. [67] use the legs for pushing and box lifting tasks by introducing additional support legs named prongs to enhance the robustness and manipulation capabilities of quadrupedal robots. Jeon et al. [32] adopted a similar hierarchical framework to us, but does not involve any visual inputs, and only learned a quadruped for pushing large objects without an additional arm.
@@
The system we design highly differs from the aforementioned works, as our work is a fully autonomous and vision-based mobile manipulation system on quadruped robots, with limited human intervention. The robot is controlled and planned by trained policies and can adapt to objects with varying heights through its whole-body behavior, and emergent retrying behaviors when the task fails. Due to our design of visual inputs, our system can work both indoors and outdoors with no external limits, as illustrated in Fig. 1, compared with almost all of the above-mentioned works that are only tested in limited lab scenarios. In Fig. 2 we show why VBC is highly advantaged of the non-whole-body methods [81, 79, 4], for the body control behavior allowing much more flexible poses (e.g., bending) to reach objects in random heights, like the one on the ground.
@@
Figure 3: Training a vision-based whole-body manipulator. Our VBC framework trains a low-level control policy and high-level planning policy end-to-end separately for whole-body manipulation.
@@ III-B Low-Level Policy for Whole-Body Goal-Reaching
Arm control. Previous work [22] has revealed that whole-body RL policy is good at controlling the end-effector position but struggles in controlling the end-effector orientation. Unfortunately, such behaviors are critical for our manipulation tasks. Thereafter, we take an off-the-self Inverse Kinematic (IK) Solver to convert the end-effector pose command into target joint angles for controlling the arm. Formally, IK computes the increment of the current joint position $\Delta\mathbf{\theta}\in\mathbb{R}^{6}$
@@ C.1. Privileged Teacher Policy
High-level Actions. The high-level policy $\pi^{\text{high}}$ outputs actions that determine the desired tracking position for the low-level policy. The high-level action is formally defined as:
@@
where $\mathbf{p}_{t}^{\text{cmd}}\in\mathbb{R}^{6}$ is the increment of gripper pose, $\mathbf{v}_{t}^{\text{cmd}}\in\mathbb{R}^{2}$ is the command of quadruped linear velocity and yaw velocity and $p_{t}^{\text{gripper}}\in\{0,1\}$ indicates whether the gripper should be close or open.
@@
C.2. Visuomotor Student Policy
We find robust and adaptive picking-up behaviors emerge from the privileged state policy training. Nonetheless, it is assumed that object poses and shape features, which are inaccessible during deployment, are taken as input. To deploy our policy in the real world, we need to use easily accessible real-world observations to infer (representations of) these properties.
Observations and actions. As camera images are easily accessible both in simulation and the real world, we use camera images that implicitly contain the information of object state and shape to train our teacher policy. To mitigate the deployment gap, we use only depth images, as RGB images in the real world tend to differ greatly from those in the simulator. Besides, we use segmented depth images instead of whole depth images to avoid the negative influence of surroundings. Formally, the observation of our vision policy is
@@ C.2. Visuomotor Student Policy
Imitation learning for student distillation. To deploy our policy in the real world, we distill our state-based high-level planning policy into a vision-based policy using DAgger. We warm up by teacher sampling the initial transition data. This is like behavior cloning (BC) to find a good initialization of the visuomotor student policy to improve the training efficiency. Then we sample transitions using the student policy and request correction from the teacher. We adopt augmentations like RandomErasing, GaussianBlur, GaussianNoise, and RandomRotation for the depth images. More details can be found in Appendix II-B.
Policy network architecture. Our vision-based high-level policy network comprises a CNN, a GRU, and a two-layer MLP that outputs the high-level actions. We concatenate the images and other states as observations. The images are passed through the two-layer CNN network with kernel size being $5$ and $3$, respectively, and then encoded as a $64$-dim latent vector. Then, we concatenate this latent vector with other states as a flattened vector. This vector is then passed to a one-layer GRU with hidden size 128, followed by a two-layer MLP with hidden size 64. We use ELU as the activation.
@@ IV Experiments
Figure 10: Success rates of different student policies at different heights, tested in the simulator. Each distinct object for more than 100 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps ($\sim$ 18 seconds). The dots represent the mean performance of all objects, the upper and the lower bound represent the maximum and the minimum success rate overall 7 categories, where we also note their category name.
@@ IV-A Simulation Results
TABLE I: Success rates of VBC compared to baseline method on seven categories of objects, tested in the simulator. Each distinct object for more than 300 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps. With the 3D shape feature, the algorithm is better at handling irregular objects. Detailed descriptions of each method can be referred to Section IV-A.
@@
Evaluation principles and baselines. In the simulation, when the environment is reset and a new trail begins, both the robot and the objects are initialized with random positions and poses, and it’s considered a success if an object is picked up and a failure when the object falls down the table or is not successful in 150 high-level steps. During experiments, we compare both the privileged teacher policies of VBC against baselines, and their visuomotor student policies over $34$ kinds of different objects: 1) VBC w.o. shape feature: a visuomotor policy that is trained the same as VBC, but does not access the pre-trained object shape features during the teacher training stage. This aims to show the necessity of the object shape features during teacher training. We only compare the teacher’s performance as the only difference comes from the teacher’s side. 2) Floating base: a floating base policy with a perfect low-level navigation ability but without whole-body behavior; in other words, the robot can always follow the given velocity commands and body height commands ranging from 0.4 to 0.55 meters (we set 0.4 meters as the lower bound as the default controller in real-world only allows 0.47 meters in minimum). This is to show the indispensability of the whole-body behavior. We compare both the teacher and its corresponding student policy using the same policy network structure. 3) Non-hierarchical: a unified policy trained with low-level policy and visuomotor high-level policy jointly in an end-to-end style. This policy takes the observation of our low-level and high-level policies together. It outputs all the target positions of 12 robot joint angles and the target pose of the gripper. We failed in training such a policy, demonstrating that the training pipeline of VBC is inevitable.
Figure 12: Visualization of robot behavior and corresponding commands during the test time. Particularly, we plot two kinds of high-level commands given by the high-level policy, as described in Section III-C, i.e., the velocity commands (including the quadruped linear velocity and yaw velocity); along with the gripper command (indicating whether the gripper should be close or open).
@@ IV-B Real-World Experiments
Evaluation principles and baselines. We deploy the trained visuomotor student policy of VBC as tested in the simulator directly into the real world. We In the real world, we test how many times the robot can pick up the assigned object in 5 trials (i.e., the robot is allowed to continuously retry four times when it fails to grasp without resetting), as we find that the robot emerges to have retrying behavior, as will be described below. We choose this as a measurement since we find that the emergent retrying behaviors improve the robustness, reliability, and success rate of finishing a job. We consider it a success when the object is picked more than 0.1m higher than the plane where it is placed and a failure when the robot tries more than five times or the object falls from its surface. Each object is evaluated over three different heights, as stated before, and five resets are permitted for each height. In this part, we compare VBC with a similar high-level policy trained with stationary legs. During the test of the real-world mobile manipulation task, we put the objects at an appropriate distance that the cameras mounted on the robot can observe, thus they can be manually annotated using our developed TrackingSAM tool. After the annotation, the robot is able to walk by and pick up the objects in a fully autonomous way.
Emergent retrying behaviors. It is worth noting that during simulation training, the robot emerges to learn a retrying behavior. In other words, when the robot fails to grasp an object, it will automatically retry to grasp without any human intervention. This shows an advantage and a clear difference compared to model-based methods.
Pickup performance. We collect all results and conclude the averaged performance over all objects on each height setting and show in Fig. 11. It is worth noting that our method allows for autonomous retrying behaviors when one pick-up trial fails. Obviously, VBC surpass the baseline method on all settings. To highlight, the baseline methods fail at 0.0m and 0.3m, as the default controller keeps a fixed robot height, similar to the floating base baseline. Even on the 0.5m setting, where the teleoperated baseline without whole-body behavior is equivalent to a static robotics arm, VBC is generally better. Besides, VBC shows great generalization ability on unseen shapes, thanks to the training strategy and observation design in the training pipeline.

### v4

@@ Abstract
We study the problem of mobile manipulation using legged robots equipped with an arm, namely legged loco-manipulation. The robot legs, while usually utilized for mobility, offer an opportunity to amplify the manipulation capabilities by conducting whole-body control. That is, the robot can control the legs and the arm at the same time to extend its workspace. We propose a framework that can conduct the whole-body control autonomously with visual observations. Our approach, namely Visual Whole-Body Control (VBC), is composed of a low-level policy using all degrees of freedom to track the body velocities along with the end-effector position, and a high-level policy proposing the velocities and end-effector position based on visual inputs. We train both levels of policies in simulation and perform Sim2Real transfer for real robot deployment. We perform extensive experiments and show significant improvements over baselines in picking up diverse objects in different configurations (heights, locations, orientations) and environments.
@@ I Introduction
An important aspect of loco-manipulation is the opportunity to conduct whole-body control using all degrees of freedom of the robot. While the legs are typically utilized for mobility, there is a large potential for controlling all the joints of the legs and arms together to amplify manipulation capabilities. We humans do this, too: We use our legs to stand up to reach a book placed on a shelf; We bend our legs to tie our shoes. Fig. 1 shows the tasks our robot can do when controlling the whole-body joints together. For example, our robot can go to the side of the road and pick up the trash on the grass. This will be unachievable without bending the front legs of the robot. The arm alone will not be long enough to reach the ground.
@@ II Related Work
Legged Locomotion. For decades, legged locomotion has been an essential topic in the field of robotics and has gained much success. Classical methods solve the legged locomotion task through model-based control to define controllers in advance. For example, Miura and Shimoyama [50], Sreenath et al. [61] utilized traditional methods to enable the walking of biped robots. MIT cheetah 3 [7], ANYmal-a [31] by ETH, and Atlas [19] by Boston Dynamics are good examples of robots designed with a robust walking controller. Based on these developed robots, researchers have been able to develop advanced controllers for even more agile locomotion tasks, such as jumping or walking over varying obstacles [53, 51, 26, 52, 3, 38]. To make a robust controller that can adapt to varying environments, researchers integrate some sensor data such as elevation maps [18, 40, 35, 6, 17] and state estimators [8, 66, 10, 74, 65]. Although classical methods have greatly succeeded in legged locomotion, generalizability in pre-designed controllers would require careful and tedious engineering efforts. With the fast development of deep neural networks, learning-based end-to-end methods have become a simple way to gain robust controllers with strong generalization abilities. Learning-based methods often train a robust controller in physics simulators such as PyBullet [14] and Isaac Gym [47] and then perform sim-to-real transfer to a real robot. For example, Rudin et al. [58] uses Isaac Gym with GPU to enable a quadruped robot to walk in only several minutes, Kumar et al. [36, 37] proposed the RMA algorithm to make the sim-to-real transfer process easier, and more robust, Lee et al. [41], Miki et al. [49], Agarwal et al. [1], Yang et al. [73], Yu et al. [80], Cheng et al. [13], Zhuang et al. [82], Fu et al. [21], Margolis and Agrawal [48], Li et al. [44] enable the legged robots to walk or run on varying terrains, and Fuchioka et al. [24], Escontrela et al. [16], Peng et al. [54], Yang et al. [71], Wang et al. [64], Li et al. [43], Kumar et al. [39] combine RL with imitation learning to produce naturalistic and robust controllers for legged robots. Kim et al. [33] map human motions to a quadruped robot to complete a wide range of locomotion and manipulation tasks.
@@
Legged Loco-Manipulation A set of works has exploited legged loco-manipulation to take advantage of legged robots. Bellicoso et al. [5] designs an online motion planning framework, together with a whole-body controller based on a hierarchical optimization algorithm to achieve loco-manipulation, Fu et al. [22] proposed an effective method to train a whole-body controller end-to-end through DRL. Ma et al. [46] combined model-based manipulator control and learning-based locomotion, but only showed the response to external push instead of achieving manipulation tasks. Notably, these works all require human involvement and teleoperation to execute and achieve tasks, rather than operating autonomously. Zhang et al. [81], Yokoyama et al. [79], Arcari et al. [4] all performed an autonomous mobile manipulation system on quadruped robots while they utilized the default model-based controller provided through the robot APIs (only control robot locomotion in a static height) and only focused on the manipulation part without any whole-body behaviors to adapt to different heights of objects. Ferrolho et al. [20] proposed a trajectory optimization framework based on a robustness metric for solving complex loco-manipulation tasks, but the task is assigned and the target positions are calculated via a motion capture system, thus only works in limit. Sleiman et al. [60] realized a bilevel search strategy for motion planning, by incorporating domain-specific rules, combing trajectory optimization, informed graph search, and sampling-based planning, enabling robots to perform complex tasks. They require the user to provide a dense description of the scene to initialize the solver. Zimmermann et al. [83] also utilized trajectory optimization to enable a legged robot to walk and pick up objects. However, they only show limited surroundings and simple objects. Besides these works with arm-based configuration, Wolfslag et al. [67] use the legs for pushing and box lifting tasks by introducing additional support legs named prongs to enhance the robustness and manipulation capabilities of quadrupedal robots. Jeon et al. [32] adopted a similar hierarchical framework to us, but does not involve any visual inputs, and only learned a quadruped for pushing large objects without an additional arm.
@@
The system we design highly differs from the aforementioned works, as our work is a fully autonomous and vision-based mobile manipulation system on quadruped robots, with limited human intervention. The robot is controlled and planned by an end-to-end policy integrated with perception and control, able to adapt to objects with varying heights through its whole-body behavior, and emerges retrying behaviors when the task fails. Due to our design of visual inputs, our system can work both indoors and outdoors with no external limits, as illustrated in Fig. 1, compared with almost all of the above-mentioned works that are only tested in limited lab scenarios. In Fig. 2 we show why VBC is highly advantaged of the non-whole-body methods [81, 79, 4], for the body control behavior allowing much more flexible poses (e.g., bending) to reach objects in random heights, like the one on the ground.
@@
Figure 3: Training a vision-based whole-body policy for loco-manipulation. Our VBC framework trains a low-level control policy and a high-level planning policy using reinforcement learning and imitation learning.
@@ III-B Low-Level Policy for Whole-Body Goal-Reaching
Arm control. Previous work [22] has revealed that whole-body RL policy is good at controlling the end-effector position but struggles in controlling the end-effector orientation. Unfortunately, such behaviors are critical for our manipulation tasks. Thereafter, we solve the inverse kinematic (IK) with the pseudoinverse Jacobian method222 More advanced IK solver can be exploited for better performance under certain circumstances, but the pseudoinverse Jacobian method works well enough for the problem studied in this paper. to convert the end-effector pose command into target joint angles for controlling the arm. Formally, IK computes the increment of the current joint position $\Delta\mathbf{\theta}\in\mathbb{R}^{6}$
@@ C.1. Privileged Teacher Policy
High-level Actions. The high-level policy $\pi^{\text{high}}$ outputs actions that determine the velocities and the target end-effector position for the low-level policy. The high-level action is formally defined as:
@@
where $\mathbf{p}_{t}^{\text{cmd}}\in\mathbb{R}^{6}$ is the increment of gripper pose, $\mathbf{v}_{t}^{\text{cmd}}\in\mathbb{R}^{2}$ is the command of quadruped linear velocity and yaw velocity and $p_{t}^{\text{gripper}}\in\{0,1\}$ indicates the gripper status (close or open).
@@
C.2. Visuomotor Student Policy We find robust and adaptive picking-up behaviors emerge from the privileged teacher policy training, with inaccessible object poses and shape features in reality. To deploy our policy in the real world, we must use easily accessible real-world observations like images.
Observations and actions. Although camera images are easily accessible both in simulation and the real world, there is a huge gap for colored images between the simulation and the real world. This persuades us to use depth images only. To further decrease the noise and gap from backgrounds, we use the segmented depth images and also their segmentation masks. Formally, the observation of our vision policy is
@@
Imitation learning for student distillation. To deploy our policy in the real world, we distill our state-based high-level planning policy into a vision-based policy using DAgger. We warm up by teacher sampling the initial transition data. This is like behavior cloning (BC) to find a good initialization of the visuomotor student policy to improve the training efficiency. Then we sample transitions using the student policy and request correction from the teacher. We find this is much more efficient and effective than directly learning a visuomotor student policy using RL. We adopt augmentations like RandomErasing, GaussianBlur, GaussianNoise, and RandomRotation for the depth images. More details can be found in Appendix II-B.
Policy network architecture. Our vision-based high-level policy network comprises a CNN and a two-layer MLP that outputs the high-level actions. We concatenate the images and other states as observations. We stack 4-step images which are passed through the two-layer CNN network with kernel size being $5$ and $3$, respectively, and then encoded as a $64$-dim latent vector. Then, we concatenate this latent vector with other states as a flattened vector. This vector is then passed to a two-layer MLP with a hidden size of 64. We use ELU as the activation.
@@ IV Experiments
Figure 10: Success rates of different student policies at different heights, tested in the simulator. Each distinct object for more than 100 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps ($\sim$ 18 seconds). The dots represent the mean performance of all objects, and the upper and the lower bound represent the maximum and the minimum success rate overall 7 categories, where we also note their category name.
@@ IV-A Simulation Results
TABLE I: Success rates of VBC compared to baseline method on seven categories of objects, tested in the simulator. Each distinct object for more than 300 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps. With the 3D shape feature, the algorithm is better at handling irregular objects. Detailed descriptions of each method are stated in Section IV-A.
@@
Evaluation principles and baselines. In the simulation, when the environment is reset and a new trail begins, both the robot and the objects are initialized with random positions and poses, and it’s considered a success if an object is picked up and a failure when the object falls down the table or is not successful in 150 high-level steps. During experiments, we compare both the privileged teacher policies of VBC against baselines, and their visuomotor student policies over $34$ kinds of different objects: 1) VBC w.o. shape feature: a visuomotor policy that is trained the same as VBC, but does not access the pre-trained object shape features during the teacher training stage. This aims to show the necessity of the object shape features during teacher training. We only compare the teacher’s performance as the only difference comes from the teacher’s side. 2) Floating base: a floating base policy with a perfect low-level navigation ability but without whole-body behavior; in other words, the robot can always follow the given velocity commands and body height commands ranging from 0.4 to 0.55 meters (we set 0.4 meters as the lower bound as the default controller in real-world only allows 0.47 meters in minimum). This is to show the indispensability of the whole-body behavior. We compare both the teacher and its corresponding student policy using the same policy network structure. 3) Non-hierarchical: a unified policy trained with low-level policy and visuomotor high-level policy jointly in an end-to-end style. This policy takes the observation of our low-level and high-level policies together. It outputs all the target positions of 12 robot joint angles and the target pose of the gripper. We failed in training such a policy, indicating the effectiveness and necessity of the training pipeline of VBC. We design these baselines to provide a valuable analysis of the design of each module and part.
Figure 12: Visualization of robot behavior and corresponding commands during the test time. Particularly, we plot the two kinds of high-level commands given by the high-level policy, as described in Section III-C, i.e., the velocity commands (including the quadruped linear velocity and yaw velocity); along with the gripper command (close or open).
@@ IV-B Real-World Experiments
Evaluation principles and baselines. We deploy the trained visuomotor student policy of VBC as tested in the simulator directly into the real world. In the real world, we test how many times the robot can pick up the assigned object in 5 trials (i.e., the robot is allowed to continuously retry four times when it fails to grasp without resetting), as we find that the robot emerges to have retrying behavior, as will be described below. We choose this as a measurement since we find that the emergent retrying behaviors improve the robustness, reliability, and success rate of finishing a job. We consider it a success when the object is picked more than 0.1m higher than the plane where it is placed and a failure when the robot tries more than five times or the object falls from its surface. Each object is evaluated over three different heights, as stated before, and five resets are permitted for each height. Regarding baselines, we find it hard to build a fair autonomous method for comparison, so we compared VBC with the default Unitree base controller (allowing the heights of the arm base to vary from 0.47m to 0.55m) combined with a similar high-level policy trained on stationary legs, in which the robot has no whole-body behaviors but also retains the retrying. We teleoperated the robot to be closed enough to the object before the policy started to grasp. During the test of the real-world experiments, we put the objects at a distance that the cameras mounted on the robot can observe, thus they can be manually annotated using our developed TrackingSAM tool. After the annotation, the robot walks by and picks up the objects in a fully autonomous way.
Emergent retrying behaviors. It is worth noting that during simulation training, the robot emerges to learn a retrying behavior. In other words, when the robot fails to grasp an object, it will automatically retry to grasp without any human intervention. This shows an advantage and a clear difference compared to model-based methods. In our observation, as long as the visual mask is on-track/not-lost, the success rates benefit from 1) the change of the object pose in the early tries; and 2) the movement of the robot body.
Pickup performance. We collect all results and conclude the averaged performance over all objects on each height setting and show in Fig. 11. It is worth noting that our method allows for autonomous retrying behaviors when one pick-up trial fails. Obviously, VBC surpasses the baseline method on all settings. To highlight, the baseline methods fail at 0.0m and 0.3m, as the default controller keeps a fixed robot height, similar to the floating base baseline. Even on the 0.5m setting, where the teleoperated baseline without whole-body behavior is equivalent to a static robotics arm, VBC is generally better. Besides, VBC shows great generalization ability on unseen shapes, thanks to the training strategy and observation design in the training pipeline.